OpenStudy (anonymous):

Does anyone know how to figure out this problem sqrt 3 (x+6=-12), and if it is an extraneous solution?

hartnn (hartnn):

can you post the problem clearly ? is it \(\sqrt{3(x+6)} = -12\) ?

hartnn (hartnn):

if so, can the answer to a square root be ever negative ?

OpenStudy (anonymous):

the 3 is outside the square root

OpenStudy (anonymous):

and yes it can be negative

hartnn (hartnn):

3 is outside ? so \(3 \sqrt {x+6} =-12\) ??

OpenStudy (tkhunny):

Or \(\sqrt[3]{x+6} = -12\)